bewellplus.gsu.edu
00Jzlyne00 00Jzlyne00
  • 26-01-2015
  • Physics
Answered

Support that P(e)=0.76,P(F)=0.45,and P(E and F)=0.32. What is P (E or F)?

Verified Answer :

AL2006
AL2006 AL2006
  • 26-01-2015


The probability of (E and F) = P(e) times P(f) = 0.76 x 0.45 = 0.342 .

The probability of (e OR f) =

                                   1 - P(not e AND not f) =

                                   1 - P(not e) x P(not f) =

                                   1 - (1 - 0.76) x (1 - 0.32) =

                                   1 - ( 0.24 )   x   ( 0.68 )  =  1 - 0.1632  =  0.8368.


Answer Link

Related Questions

Market pressures tend to move prices toward an
You use the length of your shoe to measure your room Your shoe is 975 inches long You walk across the room heel to toe and count 175 shoe lengthsAbout how long
Watch is a food chain
Robert is writing a paper for language arts he can write 23 of a page in 14 of an hour at what rate is robert writing
How did Europeans affect Africa in the 1800s How did Africans respond in the 1900s
when scanning the front page of a newspaper we often check the
Corey bought 112 feet of porch boards at the lumberyard If each foot cost 159how much change would he get back from 200 A 88 B 17808 C 2192 D 1379 Show the work
Which of these is responsible for ridding the body of waste
A right triangle has three sides with the same length two sides that are parallel at least two congruent sides two sides that are perpendicular
Whats 63 percent of fifty nine